OpenStudy (anonymous):

Use an appropriate Riemann sum to evaluate the limit lim((1^5 + 2^5 + 3^5 + ... + n^5)/ (n^6)) as n -> infinity

OpenStudy (anonymous):

So it would be? \[\lim_{n \rightarrow \infty}(1^5/n^6) + \lim_{n \rightarrow \infty}(2^5/n^6) +...\lim_{n \rightarrow \infty}(n^5/n^6)\]

OpenStudy (anonymous):

Wouldn't that equal 0?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

Yeah scratch that...Just realized something

OpenStudy (anonymous):

Yeah I think I'm supposed to evaluate it using right endpoints, i.e. the rectangle method, but I'm not sure how to do that in a limit where n->infinity

OpenStudy (freckles):

Yeah scratch that...Just realized something\[\lim_{n \rightarrow \infty} \frac{1}{n^6}\sum_{i=1}^{n}i^5=\lim_{n \rightarrow \infty}\frac{1}{n^6} \cdot \frac{1}{12}n^2(n+1)^2(2n^2+2n-1)\]

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}\frac{2n^6 +\cdots }{12n^6 +\cdots }\]

OpenStudy (freckles):

you only care about the coefficient of the term with highest exponent on top and bottom

OpenStudy (freckles):

the degree of the top is the same as the deg on the bottom the deg is actually 6 on top and bottom so we only need to look at 2/12

OpenStudy (anonymous):

How did you find that \[\sum_{i=1}^{n}i^5 = 1/12(n^2(n+1)^2(2n^2+2n-1)\] I remember reading that in my textbook but I can't find it again. What is this called?

OpenStudy (freckles):

your question says to use an appropriate riemann sum

OpenStudy (freckles):

that is what that is

OpenStudy (anonymous):

I see it now. Alright, so the answer should be 1/6?

OpenStudy (freckles):

yes 2/12 reduces to 1/6

OpenStudy (anonymous):

Thank you so much

OpenStudy (freckles):

np